OpenStudy (anonymous):

Describe the nature of the roots of this equation. 2x^2-x+1=0 A. Two real, irrational roots B. One real, double root C. Two complex roots D. Two real, rational roots

OpenStudy (anonymous):

well, did u find the roots?

OpenStudy (anonymous):

nah, how did u get that?

OpenStudy (cwrw238):

discriminant b^2 - 4ac = (-1)^2 - 4*2* 1 = 1 - 8 = -7 the square root of a negative are imaginary there are 2 complex roots

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

thank you! @cwrw238 :)

OpenStudy (cwrw238):

yw